One endpoint P and the midpoint M of line segment PQ are given. Find the coordinates of the other endpoint,Q.
A. P (1,4);M (3,5)

Answer by Cromlix(4381)

Hi there,
PM/MQ = 1/1
m - p/q - m = 1/1
Cross multiply.
m - p = q - m
q = 2m - p
q = 2(3) - 1
.........(5) - 4
Q = (5)
........(6)
Hope this helps :-)